OpenStudy (anonymous):

Find the critical numbers of f(x) = x^(4/5)(7x-28)^2

OpenStudy (anonymous):

\[f(x)=x^{\frac{4}{5}}(7x-28)^2\] right?

OpenStudy (anonymous):

Yep! :)

OpenStudy (anonymous):

I know I have to find where the derivative = 0 or does not exist... Things just get a little bit messy when I derive it

OpenStudy (anonymous):

it is a huge mess, that is why

OpenStudy (anonymous):

Lol XD

OpenStudy (anonymous):

one place it will not exist is at \(x=0\) since the derivative of \(x^{\frac{4}{5}}\) is \[\frac{4}{5}x^{-\frac{1}{5}}=\frac{4}{5}\frac{1}{\sqrt[5]{x}}\]

OpenStudy (anonymous):

Ohh right!

OpenStudy (anonymous):

you gotta use the chain rule and product rule to find the derivative or else expand the whole thing out and use the power rule repeatedly you still have to add up the fractions in any case

OpenStudy (anonymous):

it is a raft of boring algebra to get the derivative as one fraction i could walk you through it probably, but best idea is to cheat

OpenStudy (anonymous):

turns out the numerator is \[98(7x^2-36x+32)\] and you can set that equal to zeros and solve via the quadratic formula

OpenStudy (anonymous):

Yep I used chain rule and product rule and I got this... \[f'(x) = \frac{ 4 }{ 5 }x ^{-1/5}(7x-28)^2+14x ^{4/5}(7x-28)\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Oh how did you get that?

OpenStudy (anonymous):

oh wait how did you get the second part?

OpenStudy (anonymous):

yeah i think that is wrong

OpenStudy (anonymous):

Isn't it product rule?

OpenStudy (anonymous):

oh no it isn't my mistake you are right

OpenStudy (anonymous):

now here is what you have to do

OpenStudy (anonymous):

in order to write it as one fraction, (since the first part is a fraction) multiply top and bottom by \(\sqrt[5]{x}\)

OpenStudy (anonymous):

Oh so we multiply it on the second piece only?

OpenStudy (anonymous):

rational exponents are your friend when it comes to taking derivatives, but not when it comes to finding critical points or actually evaluating the derivative you need to get it out of exponential notation and write it as radicals

OpenStudy (anonymous):

\[\frac{4(7x-28)^2}{5\sqrt[5]{x}}+14\sqrt[5]{x^2}(7x-28)\] is what you are looking at

OpenStudy (anonymous):

to put it over the same denominator multiply the second part top and bottom by \(\sqrt[5]{x}\) which is nice because the radical will go away

OpenStudy (anonymous):

the numerator will be \[4(7x-28)^2+14x(7x-28)\] which you can do the algebra with to get a quadratic to solve

OpenStudy (anonymous):

Ahh I see!! So all I have to do now, is expand the numerator and set it = 0 and those are gonna be my critical numbers right?

OpenStudy (anonymous):

damn i made a typo, but you can probably spot it

OpenStudy (anonymous):

Hmmm... where?

jimthompson5910 (jim_thompson5910):

This is what I'm getting \[\large f(x) = x^{4/5}(7x-28)^2\] \[\large f \ '(x) = \frac{4(7x-28)^2}{5x^{1/5}}+14x^{4/5}(7x-28)\] \[\large f \ '(x) = \frac{4(7x-28)^2}{5x^{1/5}}+14x^{4/5}(7x-28) {\color{red}{\times \frac{5x^{1/5}}{5x^{1/5}}}}\] \[\large f \ '(x) = \frac{4(7x-28)^2}{5x^{1/5}}+\frac{70x(7x-28)}{5x^{1/5}}\] \[\large f \ '(x) = \frac{4(7x-28)^2+70x(7x-28)}{5x^{1/5}}\] \[\large f \ '(x) = \frac{98(x-4)(7x-8)}{5x^{1/5}}\]

OpenStudy (anonymous):

Oh the 5 was missing!! Thanks so much @jim_thompson5910 ! :)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

Expanding it, I get huge numbers, is that how it's supposed to be??

jimthompson5910 (jim_thompson5910):

set it equal to 0 and solve for x

jimthompson5910 (jim_thompson5910):

namely the numerator

OpenStudy (anonymous):

How did you get the last line @jim_thompson5910 ?

OpenStudy (anonymous):

lol i was missing the five too doh!

OpenStudy (anonymous):

also os has be acting up, i kept trying to send the wolfram line http://www.wolframalpha.com/input/?i=x^%284%2F5%29%287x-28%29^2

jimthompson5910 (jim_thompson5910):

4(7x-28)^2 + 70x(7x-28) 2(7x-28)*[2(7x-28) + 35x] ... factor out 2(7x-28) 2(7x-28)*(14x-56 + 35x) 2(7x-28)*(49x - 56) 2[7(x-4)]*[7(7x-8)] 98(x-4)(7x-8)

OpenStudy (anonymous):

Ohhh ok! Yeah I was confused on how to simplify that! XD Geee lol

OpenStudy (anonymous):

Then I have to simplify the 2 parentheses right?

jimthompson5910 (jim_thompson5910):

yeah set it equal to 0 and use the zero product property

OpenStudy (anonymous):

Ok im gonna give that a try! :D

OpenStudy (anonymous):

Thanks so much @jim_thompson5910 and @satellite73 !!! :D

jimthompson5910 (jim_thompson5910):

you're welcome